# Fairness in AI-Driven Recruitment: Challenges, Metrics, Methods, and Future Directions

### v2

@@ Abstract
The recruitment process is crucial to an organization’s ability to position itself for success, from finding qualified and well-fitting job candidates to impacting its output and culture. Therefore, over the past century, human resources experts and industrial-organizational psychologists have established hiring practices such as attracting candidates with job ads, gauging a candidate’s skills with assessments, and using interview questions to assess organizational fit. However, the advent of big data and machine learning has led to a rapid transformation in the traditional recruitment process as many organizations have moved to using artificial intelligence (AI). Given the prevalence of AI-based recruitment, there is growing concern that human biases may carry over to decisions made by these systems, which can amplify the effect through systematic application. Empirical studies have identified prevalent biases in candidate ranking software and chatbot interactions, catalyzing a growing body of research dedicated to AI fairness over the last decade. This paper provides a comprehensive overview of this emerging field by discussing the types of biases encountered in AI-driven recruitment, exploring various fairness metrics and mitigation methods, and examining tools for auditing these systems. We highlight current challenges and outline future directions for developing fair AI recruitment applications, ensuring equitable candidate treatment and enhancing organizational outcomes.
Index Terms:
@@
Artificial intelligence, bias, ethical hiring, fairness, recruitment
I Introduction
The recruitment process is crucial for an organization’s success, influencing both the selection of qualified candidates and the overall organizational culture and performance. Traditional hiring practices, such as job advertisements, skill assessments, and personality tests [1], have been extensively studied by human resource (HR) experts and industrial-organizational (I-O) psychologists. These practices significantly impact organizational outcomes, including performance, employee turnover, and productivity [2, 3]. Furthermore, these outcomes affect millions of job applicants each year, given that the median employee tenure in the United States is 4.2 years and individuals hold an average of 12.3 jobs over their careers [4]. In response to evolving needs, organizations are increasingly adopting innovative recruitment strategies to enhance long-term success and optimize the process of hiring the right jobseekers.
Over the past thirty years, the hiring process has experienced a significant transformation, progressively shifting towards digital recruiting. This evolution can be categorized into three distinct phases. The first phase, digital recruiting 1.0, emerged in the mid-to-late 1990s, marked by the advent of online job descriptions and the ability for jobseekers to submit electronic resumes and applications [5]. The second phase, digital recruiting 2.0, began in the following decade with the rise of job board sites like Indeed and Monster [5]. These platforms aggregated job postings from various organizations, offering jobseekers search and recommendation features to identify positions aligning with their interests and career goals [5]. The current phase, digital recruiting 3.0, has been driven by the integration of artificial intelligence (AI), which has revolutionized recruitment practices across job boards, online networking platforms, and HR applications [5]. Today, AI is extensively utilized by organizations for diverse functions, including job advertisement creation, applicant tracking systems, conducting video and phone interviews, and implementing gamified assessments [6, 1, 5].
I-A Motivation
@@
The growing reliance on AI in recruitment is driven by its demonstrated effectiveness in various HR tasks. A key area is job applicant screening, where digital methods have been shown to expedite the process while potentially reducing human bias [6]. For large organizations handling thousands or even millions of applications, automated recruitment methods have become increasingly essential [5, 7, 6]. AI also plays a significant role in crafting job advertisements, with large language models (LLMs) like ChatGPT being used to draft and outline the necessary skills and qualifications, thereby attracting suitable candidates [8]. Additionally, LLMs are employed to generate interview questions and refine communications with job candidates [8]. Given the significant shift towards remote work, particularly following the COVID-19 pandemic, the reliance on AI for hiring and recruitment decisions is expected to increase substantially [9, 10]. This integration of AI into recruitment processes is underscored by recent statistics showing that, as of 2019, 88% of organizations globally utilized some form of AI in their recruitment activities. Specifically, 41% used AI-based chatbots for candidate engagement, 44% used AI for candidate identification from social media and other public data, and 43% used AI for recommending training to candidates [11]. With the increasing availability of cloud storage and computing resources, the prevalence of AI applications in recruitment is expected to continue growing.
@@
However, as AI adoption becomes more widespread, concerns are growing that decisions made by these systems could be influenced by the biases of organizational personnel or model developers, as evidenced by several recent incidents [1]. For instance, in 2015, Google’s job recommendation system exhibited gender bias by displaying high-income job postings more frequently to men than to women [12]. Similarly, in 2017, Amazon’s AI-based candidate evaluation tool was discontinued due to its discriminatory practice of assigning lower scores to women’s resumes, attributed to the underrepresentation of female applicants in the model’s training data [13, 1]. Moreover, in 2019, Facebook’s housing and job ad delivery system faced similar issues, with job ads being skewed based on users’ gender and race [14]. These examples illustrate how biases commonly found in the hiring process, such as hiring discrimination [15], can easily transfer to AI-based systems through the data used for training algorithms. A timeline of well-known instances of bias in AI related to recruitment applications is shown in Figure 1. Therefore, AI bias in HR applications is already a significant problem, and as AI use continues to grow rapidly, its importance will escalate dramatically in the future.
I-B Related Work and Our Contributions
@@
This paper provides a comprehensive review of fairness in machine learning (ML), deep learning, and AI-based systems specifically for recruitment and hiring, building upon our previous survey [1]. Past studies and surveys have explored biases in non-AI-based recruitment, identifying issues such as confirmation bias (making decisions within the first few minutes of meeting a candidate) and expectation anchor bias (being influenced by a single piece of information), as well as examining how and where AI can be applied in recruitment [16, 17, 18, 19, 20, 21, 22]. Additionally, past research has reviewed specific components of the hiring process, such as AI-driven assessments [23], automated video interviews [24], and their perceptions by job candidates [22].
@@
However, there remains a need for a systematic review of the literature on algorithmic fairness in AI-based recruitment, encompassing all stages of the hiring process where AI can be applied. This need is underscored by the increasing prevalence of AI in recruitment, as evidenced by a 2022 survey revealing that up to 86% of employers reported using technology-driven job interviews, with many adopting automated video interviews as a preferred method [25]. Further examples of AI utilization include using LLMs to craft inclusive job descriptions, employing chatbot systems to engage candidates prior to direct recruiter interaction, and integrating AI into applicant tracking systems for resume screening [8, 6, 1, 5]. Additionally, a 2024 survey indicated that one in four recruiters were already utilizing AI in some capacity [26].
@@
This paper addresses the gap in literature by providing a systematic review of algorithmic fairness in AI-based recruitment, making several key contributions. First, we explore the definitions of fairness, trust, and justice, drawing from I-O psychology, HR practices, and legislative frameworks to provide a foundational understanding for analyzing bias in AI-driven recruitment (Section II). Second, we discuss the ethical and legal principles governing fairness in AI, highlighting current regulations and guidelines that inform fair AI practices in recruitment (Section II-B). Third, we identify and categorize various causes of bias in AI systems, including biases in training data, label definitions, feature selection, proxies, and masking (Section II-C). Our examination extends across all stages of the recruitment process—sourcing, screening, interviewing, and selection—demonstrating how bias can be introduced and mitigated at each step (Section III). Fourth, we review diverse fairness metrics, including individual fairness, group fairness, procedural fairness, and fairness in regression, discussing their relevance for employers, jobseekers, and recruitment platforms (Section IV-A). Fifth, we categorize and evaluate bias mitigation strategies into pre-processing, in-processing, and post-processing methods, highlighting their significance and application in ensuring fair AI-driven recruitment practices (Section IV-B). Sixth, we cover the limitations of existing tools and methods for bias detection and mitigation (Section IV), and outline future challenges in this domain (Section IV-C). Finally, Section V presents concluding remarks and summarizes the key findings. By addressing these critical aspects, our paper aims to set a comprehensive foundation for advancing equitable AI practices in recruitment.
II Background
@@
Before measuring bias, it is essential to define the concept of “fairness” to ensure consistent comparison and evaluation of AI in hiring. The following sections provide definitions of fairness within the context of organizational hiring (Section II-A), outline relevant legislation defining fairness (Section II-B), and discuss common causes of bias in AI systems (Section II-C).
Figure 1: Timeline of reports identifying bias in AI systems related to or influencing current recruitment applications [13, 27, 12, 28, 29, 14, 30, 31, 32, 33, 34, 35, 36, 37].
II-A Relating Fairness to Trust and Justice
@@
The terms “fairness,” “trust,” and “justice” are closely related and often used interchangeably [38]. Before defining fairness, it is useful to examine its relationship with trust and justice, particularly in the context of AI-based hiring.
@@
First, trust is defined in I-O psychology as a psychological state of belief regarding how a system, person, or organization will act in the future based on past interactions [39, 38]. Trust takes time to build and becomes more significant when perceived uncertainty and risk increase (e.g., when the risk of job loss is high) [38]. However, trust can be easily lost, and once lost, it is very difficult to rebuild, leading workers and employees to question the fairness of organizational actions [38]. Therefore, trust in AI-based hiring significantly influences a jobseeker’s perception of the fairness of an organization’s recruitment process.
@@
Trust in technology comprises human characteristics (e.g., personality and ability), environmental characteristics (e.g., cultural background and institutional factors), and technological characteristics (e.g., performance, such as reviews from other users, and goal congruence, ensuring the AI’s goals align with human/organizational goals) [40]. Specific factors that impede trust in AI include its role in technological displacement, lack of transparency, and privacy concerns arising from the collection and reliance on large training datasets [40]. A 2019 poll highlights the lack of trust in AI, revealing that 88% of Americans are skeptical of AI-driven recruitment [41, 42], and as of 2023, 71% opposed AI making final hiring decisions [43]. Therefore, even if model developers consider fairness metrics and precautions, a jobseeker’s trust will significantly affect the perceived fairness of any AI-based decisions.
@@
Organizational justice, in this context, refers to the just treatment in organizational processes and outcomes, such as recruitment [38]. Organizational justice is crucial because it significantly influences perceptions of fairness [38]. The literature identifies three primary types of organizational justice: distributive, procedural, and interactional. Distributive justice pertains to the perceived fairness of the allocation of outcomes or rewards to individuals within an organization. This can be based on the merit or equity norm (i.e., those who work harder or produce more should receive greater rewards), the need norm (i.e., rewards are distributed in proportion to the needs of individuals), or the equality norm (i.e., all individuals receive equal rewards, regardless of effort or need) [38, 44]. Procedural justice is the perceived fairness of the processes by which rewards are distributed [38, 44]. Lastly, interactional justice concerns the perceived extent to which an organization treats an individual or employee with respect. Interactional justice has two aspects: interpersonal justice, which pertains to the extent of respect and politeness accorded to individuals, and informational justice, which reflects how well an organization explains its procedures and outcomes to individuals, ensuring that communication is candid, thorough, and timely [38].
@@
II-B Fairness According to Legislation and Ethical Principles
@@
In the ML community, the FAT/ML (fairness, accountability, and transparency) group has established five guiding principles to aid developers and product managers in designing and implementing accountable automated decision-making algorithms: responsibility, explainability, accuracy, auditability, and fairness [45]. The group also recommends adherence to OECD privacy principles and the ethical principles outlined in the Belmont Report for human subjects research to ensure algorithmic accountability [46].
@@
Additionally, in 2020, the U.S. Department of Defense defined five ethical principles for its AI systems: responsible, equitable, traceable, reliable, and governable [47, 48]. Following this, numerous companies, including Google [49] and OpenAI [50], have established similar guidelines to mandate the ethical development and integration of their AI technologies.
@@
Similarly, legislation regulating technology has emerged to enhance fairness in AI systems. For instance, in 2018, the European Union enforced the General Data Protection Regulation (GDPR) to protect individuals concerning the privacy, processing, and free movement of personal data [51]. Moreover, in 2021, the European Commission established the first legal framework on AI, categorizing AI systems into different levels of risk [52]. These levels include minimal risk, such as AI-enabled video games or spam filters that pose minimal or no risk to a citizen’s rights or safety; limited risk, where AI systems like chatbots inform users they are interacting with a machine and enable them to make informed decisions before continuing use; high-risk, covering AI systems critical to infrastructure, vocational training, and essential public services like credit scoring, law enforcement, and migration; and unacceptable risk, involving AI systems that present a clear threat to the safety and rights of people [52].
@@
The National Institute of Standards and Technology (NIST) and the Federal Trade Commission (FTC) have also emphasized the need to identify and manage bias in AI [53, 54]. Specifically, the FTC has warned against the sale of racially biased systems that could prevent individuals from obtaining employment, insurance, or other benefits. The FTC highlighted three laws for AI developers and users: Section 5 of the FTC Act, which prohibits unfair or deceptive practices; the Fair Credit Reporting Act, which ensures algorithms do not unfairly deny individuals housing or credit; and the Equal Credit Opportunity Act, which prohibits algorithms from making biased credit decisions based on race, color, religion, national origin, age, sex, marital status, or public assistance status [53, 55].
@@
The Biden Administration launched the National AI Research Resource Task Force [56], as directed by Congress in the National AI Initiative Act of 2020, to advance AI research and explore AI’s implications [57]. Following this, additional legislation such as the Algorithmic Fairness Act of 2020 [58] emerged, and in 2021 alone, 17 states introduced AI regulations [59]. Notably, New York City enacted a law (effective in 2023) prohibiting the use of automated decision-making tools to screen job candidates without prior bias audits and transparency regarding their use [60]. As of 2023, 25 U.S. states have introduced AI bills, and 18 states have adopted or enacted AI legislation [61].
@@
Legislation specific to fairness in the employment and hiring context is enforced by the U.S. Equal Employment Opportunity Commission (EEOC). These laws make it illegal to discriminate against individuals based on race, color, religion, national origin, sex, sexual orientation, or gender identity or expression, as stipulated in Title VII of the Civil Rights Act of 1964, amended by The Civil Rights Act of 1991. Additionally, the Pregnancy Discrimination Act amendment to Title VII protects women against discrimination due to pregnancy or childbirth. The Equal Pay Act of 1963 prohibits wage and benefits discrimination based on sex. The Age Discrimination in Employment Act of 1967 protects individuals 40 years of age or older against age discrimination. The Americans with Disabilities Act of 1990 (Title I) and Sections 501 and 505 of the Rehabilitation Act of 1973 safeguard qualified individuals with disabilities. The Genetic Information Nondiscrimination Act of 2008 prohibits discrimination based on genetic information [62]. Furthermore, the Fourteenth Amendment to the U.S. Constitution provides due process protections to workers [63]. In 2021, the EEOC launched an initiative to ensure that AI and tools used in hiring and employment comply with these laws [64], and as of 2023, it has provided further guidelines for employers [65].
@@
Previously, for employee selection, the EEOC established a requirement to ensure the fair treatment of job candidates by preventing adverse impact due to hiring decisions [66, 67]. According to EEOC guidelines, adverse impact is determined by applying the four-fifths or eighty percent rule. This rule states that the selection rate for a protected group must be at least four-fifths (or 80%) of the rate for the group with the highest selection rate [67]. For instance, as illustrated in Table I, if the selection rate for black applicants is 50% of the selection rate for white applicants, it falls below the 80% threshold, indicating an adverse impact [67].

### v3

@@ Abstract
The recruitment process significantly impacts an organization’s performance, productivity, and culture. Traditionally, human resource experts and industrial-organizational psychologists have developed systematic hiring methods, including job advertising, candidate skill assessments, and structured interviews to ensure candidate-organization fit. Recently, recruitment practices have shifted dramatically toward artificial intelligence (AI)-based methods, driven by the need to efficiently manage large applicant pools. However, reliance on AI raises concerns about the amplification and propagation of human biases embedded within hiring algorithms, as empirically demonstrated by biases in candidate ranking systems and automated interview assessments. Consequently, algorithmic fairness has emerged as a critical consideration in AI-driven recruitment, aimed at rigorously addressing and mitigating these biases. This paper systematically reviews biases identified in AI-driven recruitment systems, categorizes fairness metrics and bias mitigation techniques, and highlights auditing approaches used in practice. We emphasize critical gaps and current limitations, proposing future directions to guide researchers and practitioners toward more equitable AI recruitment practices, promoting fair candidate treatment and enhancing organizational outcomes.
{IEEEImpStatement}
Artificial intelligence is increasingly used in recruitment for tasks such as preparing job advertisements, screening resumes, and interviewing candidates. These AI tools improve organizational efficiency and accuracy and reduce human subjectivity; however, unintended algorithmic biases can systematically disadvantage qualified applicants, reinforcing harmful stereotypes, undermining fairness, and reducing transparency in hiring decisions. This paper systematically reviews these biases, examines established methods to detect and mitigate unfairness, and highlights practical auditing approaches. By raising awareness of AI recruitment biases and how they can be addressed, this research supports organizations in improving candidate selection, enhancing employee satisfaction and retention, boosting productivity and economic performance, and strengthening public trust in AI-driven decision-making processes.
{IEEEkeywords}
@@
Algorithmic bias, algorithmic fairness, artificial intelligence, ethical hiring, human resource management, recruitment
1 Introduction
[…]
The recruitment process is crucial for organizational success, influencing employee selection, organizational culture, and overall productivity [1, 2]. Traditionally, hiring practices—such as job advertisements, skill assessments, and personality tests—have been extensively studied by human resource (HR) professionals and industrial-organizational (I-O) psychologists [3]. Recently, the recruitment landscape has undergone a significant shift toward digital methods, evolving through three phases: from online applications and digital resumes in the 1990s (digital recruiting 1.0), to centralized job aggregation platforms in the 2000s (digital recruiting 2.0), and currently, to extensive integration of artificial intelligence (AI), termed digital recruiting 3.0 [4]. Today, AI-driven tools are increasingly adopted for diverse recruitment tasks including job advertisement creation, candidate screening, video interviewing, and automated candidate assessments [5, 3, 4].
@@ 1 Introduction
The increasing reliance on AI in recruitment is driven by its demonstrated effectiveness in various HR tasks. A notable application area is job applicant screening, where digital methods can expedite the hiring process and potentially reduce human biases [5]. Large organizations handling substantial numbers of applications increasingly find automated recruitment methods essential for operational efficiency [4, 6, 5]. AI has also begun to play a significant role in crafting job advertisements, with large language models (LLMs), such as ChatGPT, used to draft outlines of necessary skills and qualifications, potentially attracting a broader range of suitable candidates [7]. Furthermore, organizations have started employing LLMs to generate interview questions and refine communications with job candidates [7]. Given the rapid shift towards remote work—accelerated by the COVID-19 pandemic—the use of AI for hiring and recruitment decisions is anticipated to further expand [8, 9]. Recent statistics indicate significant growth, with a 2019 industry survey reporting that 88% of organizations globally have experimented with AI in recruitment activities; among these, 41% employed AI-based chatbots for candidate engagement, 44% utilized AI for identifying candidates through social media and public data, and 43% leveraged AI for training recommendations [10]. With expanding access to cloud computing resources, AI integration in recruitment practices is likely to continue growing, underscoring the need to critically assess fairness and transparency of these methods.
@@
However, as AI adoption becomes more widespread, concerns are growing that decisions made by these systems could be influenced by the biases of organizational personnel or model developers, as evidenced by several recent incidents [3]. For instance, in 2015, Google’s job recommendation system exhibited gender bias by displaying high-income job postings more frequently to men than to women [11]. Similarly, in 2017, Amazon’s AI-based candidate evaluation tool was discontinued because it consistently assigned lower scores to women’s resumes because the AI model had learned gender biases present in historical hiring data, systematically favoring terms common in men’s resumes and penalizing those associated with women [12, 13]. Moreover, in 2019, Facebook’s housing and job ad delivery system faced similar issues, with job ads being skewed based on users’ gender and race [14]. These examples illustrate how biases commonly found in the hiring process, such as hiring discrimination [15], can easily transfer to AI-based systems through the data used for training algorithms. A timeline of well-known instances of bias in AI related to recruitment applications is shown in Figure 1. Therefore, AI bias in HR applications is already a significant problem, and as AI use continues to grow rapidly, its importance will escalate dramatically in the future.
Related Work and Our Contributions:
@@
This paper provides a comprehensive review of fairness in machine learning (ML), deep learning, and AI-based systems specifically for recruitment and hiring, building upon our previous survey [3]. Past studies and surveys have explored biases in non-AI-based recruitment, identifying issues such as confirmation bias (making decisions within the first few minutes of meeting a candidate) and expectation anchor bias (being influenced by a single piece of information), as well as examining how and where AI can be applied in recruitment [16, 17, 18, 19, 20, 21, 22]. Additionally, past research has reviewed specific components of the hiring process, such as AI-driven assessments [23], automated video interviews [24], and their perceptions by job candidates [22]. Furthermore, recent studies have significantly advanced our understanding of ethical and organizational aspects of AI recruitment, including sociotechnical frameworks for AI adoption in organizations [25], methods for intelligent talent selection [26], challenges of international HR management with AI integration [27], AI capability frameworks in HR management [28], and ethical concerns regarding potential dehumanization through AI algorithms in hiring processes [29].
@@
However, there remains a need for a systematic review of the literature on algorithmic fairness in AI-based recruitment, encompassing all stages of the hiring process where AI can be applied. This need is underscored by the increasing prevalence of AI in recruitment, such as using LLMs to craft inclusive job descriptions, employing chatbot systems to engage candidates prior to direct recruiter interaction, and integrating AI into applicant tracking systems for resume screening [7, 5, 3, 4]. The prevalence is further underscored by a 2022 survey revealing that up to 86% of employers reported using technology-driven job interviews, with many adopting automated video interviews as a preferred method [30].
@@
This paper addresses the aforementioned critical gap in the literature by providing a systematic review of algorithmic fairness in AI-based recruitment. It targets AI hiring practitioners and makes several key contributions. First, we explore the definitions of fairness, trust, and justice, drawing from I-O psychology, HR practices, and legislative frameworks to provide a foundational understanding for analyzing bias in AI-driven recruitment (Section 2). Second, we discuss the ethical and legal principles governing fairness in AI, highlighting current regulations and guidelines that inform fair AI practices in recruitment (Section 2.2). Third, we identify and categorize various causes of bias in AI systems, including biases in training data, label definitions, feature selection, proxies, and masking (Section 2.3). Our examination extends across all stages of the recruitment process—sourcing, screening, interviewing, and selection—demonstrating how bias can be introduced and mitigated at each step (Section 3). Fourth, we review diverse fairness metrics, including individual fairness, group fairness, procedural fairness, and fairness in regression, discussing their relevance for employers, jobseekers, and recruitment platforms (Section 4.1). Fifth, we categorize and evaluate bias mitigation strategies into pre-processing, in-processing, and post-processing methods, highlighting their significance and application in ensuring fair AI-driven recruitment practices (Section 4.2). Sixth, we cover the limitations of existing tools and methods for bias detection and mitigation (Section 4), and outline future challenges in this domain (Section 5). Finally, Section 6 presents concluding remarks and summarizes the key findings. By addressing these critical aspects, our paper aims to set a comprehensive foundation for advancing equitable AI practices in recruitment. However, we acknowledge that defining and measuring fairness in recruitment involves complexities due to competing fairness metrics and differing interpretations of employment discrimination law, which will be discussed in subsequent sections.
Scope of Review: In this survey, we collected relevant articles published by December 2024 through searches on Google Scholar, IEEE Xplore, ACM, ACL, and other publishers. We combined the keywords “AI” or “ML” with terms such as “Fairness,” “Recruitment,” “Recruitment Bias,” and specific terms representing each recruitment stage: “Candidate Sourcing,” “Candidate Screening,” “Candidate Interviews,” and “Offer Negotiation.” The following criteria are used to determine the relevance and/or irrelevance of an article to our survey:
  • Influential or foundational methods or studies in recruitment influencing AI development or understanding fairness in AI.
  • Work that studies bias, proposes a new method for studying bias, or is applied to mitigate bias in any stage of the recruitment process.
  • Any external studies and/or related works that are not directly part of recruitment, but can influence AI being used in one of the stages. For example, models such as GPT were not initially created for recruitment, but have heavily influenced the workplace [7].
2 Background: Defining Fairness
@@
Before measuring bias, it is essential to define the concept of “fairness” to ensure consistent comparison and evaluation of AI for hiring. The following sections provide definitions of fairness within the context of organizational hiring (Section II-A), relevant legislation (Section II-B), and common causes of bias in AI systems (Section II-C).
Figure 1: Timeline of notable incidents highlighting biases identified in AI systems relevant to or influencing recruitment applications [13, 31, 11, 32, 33, 14, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
2.1 Relating Fairness to Trust and Justice
@@
In organizational hiring contexts, fairness cannot be adequately understood without considering its close relationship with trust and justice. A clear conceptual connection among these terms is essential, as misunderstanding this relationship can significantly undermine both the perceived and actual fairness of AI-driven recruitment practices. The terms “fairness,” “trust,” and “justice” are closely related and often used interchangeably [47]. Before defining fairness, it is useful to examine its relationship with trust and justice, particularly in the context of AI-based hiring.
@@
First, trust, as defined in I-O psychology, is a psychological state characterized by the willingness of an individual to accept vulnerability based on positive expectations regarding the intentions, behavior, and outcomes associated with another entity—whether a person, organization, or technology. It encompasses reliability (consistent performance), integrity (adherence to ethical principles), competence (ability to perform tasks), and benevolence (consideration for others’ welfare) [48, 49, 47, 50, 51]. Rather than simply being derived from past interactions, trust also involves contextual factors, perceived risk, regulatory oversight, and individual predispositions towards uncertainty, sometimes described as a “leap of faith” [50]. Trust takes time to build and becomes more significant when perceived uncertainty and risk increase (e.g., when the risk of job loss is high) [47]. However, trust can be easily lost, and once lost, it is very difficult to rebuild, leading workers and employees to question the fairness of organizational actions [47]. Therefore, trust in AI-based hiring significantly influences a jobseeker’s perception of the fairness of an organization’s recruitment process.
@@
Trust in AI-driven recruitment technology is influenced by the characteristics of the end-user (e.g., the jobseeker’s or recruiter’s personality and ability), the organizational environment (e.g., institutional policies and workplace culture), and the recruitment system itself (e.g., accuracy, performance, transparency, user reviews, and goal congruence—the alignment between the system’s objectives and organizational or individual user goals) [52]. Specific factors that impede trust in AI include its role in technological displacement, lack of transparency, biased or unfair decisions, lack of accountability for those responsible for AI decisions, and privacy concerns arising from the collection and reliance on large training datasets [52, 53, 51]; for a detailed review of trust in AI, we direct readers to the paper by Afroogh et al. [53]. A 2019 poll highlights the lack of trust in AI, revealing that 88% of Americans are skeptical of AI-driven recruitment [54, 55], and as of 2023, 71% opposed AI making final hiring decisions [56]. Another study of recruiters reported they were cautious of the AI’s ability to make fair and accurate hiring decisions, and often found job candidates were unwilling to engage with AI assessments, potentially losing highly skilled candidates [57]. Therefore, even if model developers consider fairness metrics and precautions, a jobseeker’s trust will significantly affect the perceived fairness of any AI-based decisions. It is also important to recognize the inherent ethical considerations and power dynamics within recruitment, as the interests of jobseekers and employers often diverge. An ethical recruitment process should balance these competing interests, rather than solely serving as a mechanism for employers to filter out less-desirable candidates.
@@
Organizational justice, in this context, refers to the just treatment in organizational processes and outcomes, such as recruitment, and is a crucial part to understanding trust in AI because it significantly influences perceptions of fairness [47]. The literature identifies three primary types of organizational justice: distributive, procedural, and interactional. Distributive justice pertains to the perceived fairness of the allocation of outcomes or rewards to individuals within an organization. This can be based on the merit or equity norm (i.e., those who work harder or produce more should receive greater rewards), the need norm (i.e., rewards are distributed in proportion to the needs of individuals), or the equality norm (i.e., all individuals receive equal rewards, regardless of effort or need) [47, 58]. Procedural justice is the perceived fairness of the processes by which rewards are distributed [47, 58]. Lastly, interactional justice concerns the perceived extent to which an organization treats an individual or employee with respect. Interactional justice has two aspects: interpersonal justice, which pertains to the extent of respect and politeness accorded to individuals, and informational justice, which reflects how well an organization explains its procedures and outcomes to individuals, ensuring that communication is candid, thorough, and timely [47].
@@
2.2 Fairness According to Legislation and Ethical Principles
@@
Considerable progress has been made in establishing standards for fairness and ethical AI. Notably, the FAT/ML (Fairness, Accountability, and Transparency in Machine Learning) community has developed five guiding principles—responsibility, explainability, accuracy, auditability, and fairness—to help developers and product managers design and implement accountable automated decision-making algorithms [59]. The group also recommends adherence to OECD privacy principles and the ethical principles outlined in the Belmont Report for human subjects research to ensure algorithmic accountability [60].
@@
Additionally, in 2020, the U.S. Department of Defense defined five ethical principles for its AI systems: responsible, equitable, traceable, reliable, and governable [61, 62]. Following this, numerous companies, including Google [63] and OpenAI [64], have established similar guidelines for ethical development and integration of their AI technologies.
@@
Similarly, legislation regulating technology has emerged to enhance fairness in AI systems. For instance, in 2018, the European Union enforced the General Data Protection Regulation (GDPR) to protect individuals concerning the privacy, processing, and free movement of personal data [65]. Moreover, in 2021, the European Commission established the first legal framework on AI, grouping AI systems into different risk levels [66, 67]; this was officially adopted in 2024 [68]. These levels include minimal risk, such as AI-enabled video games or spam filters that pose minimal or no risk to a citizen’s rights or safety; limited risk, where AI systems like chatbots inform users they are interacting with a machine and enable them to make informed decisions before continuing use; high-risk, covering AI systems critical to infrastructure, vocational training, and essential public services like credit scoring, law enforcement, and migration; and unacceptable risk, involving AI systems that present a clear threat to the safety and rights of people [66]. Similar legislative considerations are being explored internationally; for example, Canada and Australia are currently developing legislation targeting fairness and employee protections in AI-driven recruitment [69, 70].
@@
The National Institute of Standards and Technology (NIST) and the Federal Trade Commission (FTC) have also emphasized the need to identify and manage bias in AI [71, 72]. Specifically, the FTC has warned against the sale of racially biased systems that could prevent individuals from obtaining employment, insurance, or other benefits. The FTC highlighted three laws for AI developers and users: Section 5 of the FTC Act, which prohibits unfair or deceptive practices; the Fair Credit Reporting Act, which ensures algorithms do not unfairly deny individuals housing or credit; and the Equal Credit Opportunity Act, which prohibits algorithms from making biased credit decisions based on race, color, religion, national origin, age, sex, marital status, or public assistance status [71].
@@
The Biden Administration launched the National AI Research Resource Task Force [73], as directed by Congress in the National AI Initiative Act of 2020, to advance AI research and explore AI’s implications [74]. Following this, additional legislation such as the Algorithmic Fairness Act of 2020 [75] and Algorithmic Accountability Act of 2022 [76] emerged, and in 2021 alone, 17 states introduced AI regulations [77]. Notably, New York City enacted a law (effective in 2023) prohibiting the use of automated decision-making tools to screen job candidates without prior bias audits and transparency regarding their use [78]. As of 2023, many U.S. states have introduced and enacted AI legislation [79].
@@
Legislation specific to fairness in the employment and hiring context is enforced by the U.S. Equal Employment Opportunity Commission (EEOC). These laws protect individuals against discrimination based on race, color, religion, national origin, sex, sexual orientation, or gender identity or expression, as stipulated in Title VII of the Civil Rights Act of 1964, amended by The Civil Rights Act of 1991. Additionally, the Pregnancy Discrimination Act amendment to Title VII protects women against discrimination due to pregnancy or childbirth. The Equal Pay Act of 1963 prohibits wage and benefits discrimination based on sex. The Age Discrimination in Employment Act of 1967 protects individuals 40 years of age or older against age discrimination. The Americans with Disabilities Act of 1990 (Title I) and Sections 501 and 505 of the Rehabilitation Act of 1973 safeguard qualified individuals with disabilities. The Genetic Information Nondiscrimination Act of 2008 prohibits discrimination based on genetic information [80]. In 2021, the EEOC launched an initiative to ensure that AI and tools used in hiring and employment comply with these laws [81], and as of 2023, it has provided further guidelines for employers [82].
@@
Previously, for employee selection, the EEOC established a requirement to ensure the fair treatment of job candidates by preventing adverse impact due to hiring decisions [83, 84]. According to EEOC guidelines, adverse impact is determined by applying the four-fifths or eighty percent rule. This rule states that the selection rate for a protected group must be at least four-fifths (or 80%) of the rate for the group with the highest selection rate [84]. For instance, as illustrated in Table 1, if the selection rate for Black applicants is 50% of the selection rate for White applicants, it falls below the 80% threshold, indicating an adverse impact [84]. From this and other legislation, two key definitions of discrimination have emerged in AI literature: disparate treatment, which refers to intentionally discriminatory practices targeted at individuals based on their protected attributes; and disparate impact, which encompasses practices that result in a disproportionately adverse effect on protected groups [84, 85, 86].